OpenStudy (anonymous):

Please help! The equation y = 2^x / 4 can be written in the form y=2^t find an expression for 't' in terms of x

OpenStudy (anonymous):

\(\frac{1}{4}=2^{-2}\) the laws of exponents

OpenStudy (anonymous):

we didnt learn exponents yet .

OpenStudy (anonymous):

\(2^{-2}\times 2^x=2^{x-2}\)

OpenStudy (anonymous):

hint: Example 1: (1,3) and (7,8) Solution: m= 8-3 7-1 = 5 6 Notice that if I had reversed the order of the points I would have gotten m= 3-8 1-7 = -5 -6 = 5 6 , but I need to be consistent, because 3-8 7-1 = 5 -6 , which is not the same thing. We can tell from this slope of 5/6 that the line goes uphill from left to right and it rises 5/6 as much as it travels from left to right. A line with slope 1 would be one that makes a 45° angle with the axes, so a line of slope 5/6 would be just a little less steep than that.

OpenStudy (anonymous):

sorry the question is \[y =\frac{2^x}{2} ,y= 2^x \]

OpenStudy (anonymous):

nw hw do v wrk this out

OpenStudy (unklerhaukus):

\[y = \frac{2^x} { 2}, \qquad y=2^t \]?

OpenStudy (anonymous):

thts the question how to express for t in terms of x

OpenStudy (anonymous):

i wrote the answer above, i don't know what else to say it is \(y=2^{x-2}\) so \(t=x-2\)

OpenStudy (anonymous):

can u please elaborate a bit more clear

OpenStudy (unklerhaukus):

\[2^t=\frac{2^x}{2}\]\[\downarrow\]\[2\times 2^t =2^x\] \[2^{t+1}=2^x\] \[t+1=x\]

OpenStudy (unklerhaukus):

\[t=x-1\]

OpenStudy (anonymous):

i dint get tht part \[2^{t+1} = 2^x\]

OpenStudy (unklerhaukus):

ok so 2 squared is really \[2^2 =2 \times 2\] 2 cubed is really \[2^3=2\times 2\times 2\] _________ so \[2^t=2\times 2 \times 2 ......\text{(t times)}\] \[2 \times 2^t =2\times (2 \times 2 ......)\text{(t times)}\] \[2 \times 2^t =2\times 2 \times 2 ......)\text{(t+1 times)}\]

OpenStudy (anonymous):

oh ok get it :)

OpenStudy (anonymous):

thnks ;)

OpenStudy (unklerhaukus):

np